# A Survey of Behavior Foundation Model: Next-Generation Whole-Body Control System of Humanoid Robots

### v3

@@ 2.2 Reinforcement Learning
TeamPlay [64], ASE [65], PHC [66], CALM [67], MTM [68],
CASE [69], InterMimic [70], MoConVQ [71], H-GAP [72],
MaskedMimic [73], HugWBC [60], HOVER [61], ModSkill [74]
@@
ICM [75], DIAYN [76], RND [77], APS [78],
ProtoRL [79], RE3 [80], ODPP [81]
@@
FB [82],FB-IL [44], FB-AWARE [45], FB-CPR [46]
Adaptation Adaptation
@@
TokenHSI [83], Belief-FB [49], Rotation-FB [49],
TaskTokens [48], ReLA [47], LoLA [47]
@@
UniHSI [84], TokenHSI [83], UniPhys [85],
CLoSD [86], LangWBC [87], LeVERB [88]
3 BFM for Humanoid Whole-body Control
In this section, we introduce representative methods for constructing and adapting BFMs towards humanoid WBC, analyzing extensively their main motivations, typical implementations, and empirical properties. Note that some approaches might have been proposed before the emergence of the concept of BFMs, yet we still discuss them as long as they adhere to the properties of BFMs or their physical meaning is analogous to BFMs.
3.1 Pre-training
Pre-training of BFMs seeks to learn reusable primitive skills and behavioral priors from large-scale data sources, creating a foundation for efficient downstream adaptation. Current approaches can be broadly categorized into three types (as depicted in Table II): goal-conditioned learning, intrinsic reward-driven learning, and the forward-backward representation learning.
3.1.1 Goal-conditioned Learning
As shown in Figure 3, goal-conditioned learning is a framework in RL where an agent’s behavior is conditioned on a specific goal or objective, typically provided as input. Unlike traditional RL, where the agent learns from raw state-action pairs without explicit task-specific guidance, goal-conditioned learning integrates the goal into the agent’s policy, enabling it to adapt its actions toward achieving that specific goal. The goal can be specified in various forms, such as a target state, an objective function, or an external task description [66, 68, 72]. The key advantage of goal-conditioned learning lies in its ability to learn a more flexible and transferable policy that can be applied to a wide range of tasks, as it directly incorporates the task’s goal during training, rather than requiring retraining for each specific task. This makes it particularly useful in environments where the agent needs to solve multiple tasks or interact with changing environments.
Skill learning from motion tracking. Among the diverse approaches to goal-conditioned learning, tracking-based learning represents a specialized form where the target behavior is explicitly defined by dense reference supervision or guidance, typically derived from motion capture data or expert demonstrations. At each time step, the agent is often trained to track the given reference motion’s joint angle or kinematic pose of the next time step [16]. The primary motivation behind tracking-based learning is that learning to track a single pose is more achievable and general than directly imitating a whole motion, especially a complex motion.
@@ 3.1.1 Goal-conditioned Learning
For example, [64] trains an agent to imitate a large amount of football motion capture data via a DeepMimic-like [16] approach, aiming to realize the complete behavior coverage for the football game. Then the agent is leveraged to sample substantial state-action pairs to train a neural probabilistic motor primitive (NPMP) model [89] and derive a low-level latent-conditioned controller. Finally, the controller is applied for further drill learning by conducting RL with a drill-specific (e.g., follow, dribble, shoot, and kick-to-target) reward function. Here, the learned low-level controller can be viewed as a BFM as it learns realistic human-like movements based on motion capture data and can be rapidly adapted to diverse higher-level drill learning.
Similarly, [65] introduces adversarial skill embeddings (ASE), a framework that learns a reusable latent space of motor skills by combining adversarial IL with unsupervised RL. Trained on unstructured motion data, ASE produces a latent-conditioned low-level controller capable of generating diverse and physically plausible behaviors, serving as a general-purpose motor prior to downstream tasks. Building on ASE, [67] proposes conditional adversarial latent models (CALM), which incorporate a conditional discriminator to enable fine-grained control over generated motions via latent manipulation. [69] further extends this line with CASE, introducing skill-conditioned IL with training techniques such as focal skill sampling and skeletal residual forces to enhance agility and motion diversity.
@@
While the methods above achieve efficient skill acquisition from large behavior datasets, [60] proposes HugWBC, which explores learning versatile locomotion skills without relying on pre-collected motion data. The framework automatically generates adaptive behaviors through a structured RL process, where a general command space dynamically produces feasible velocity, gait, and posture targets during training. By reformulating WBC as a self-supervised command-tracking problem, the work establishes a new direction for developing general-purpose humanoid controllers that learn robust skills through environmental interaction rather than data imitation.
@@
Moving beyond body-level skill learning, ModSkill [74] introduces a modular framework that decouples full-body motion into part-specific skills for individual body parts. This modularization allows for efficient and scalable learning, as each body part is controlled independently by a low-level controller driven by part-specific skill embeddings. ModSkill’s ability to focus on body-part-level skills makes it a powerful system for controlling complex motions and adapting learned behaviors across different tasks. By utilizing a skill modularization attention layer, ModSkill enhances the generalization of motor skills across various tasks like reaching or striking, further improving task-specific adaptation.
@@
From primitive skills to high-level goal execution. The success of BFMs in learning diverse primitive skills has propelled the development of more advanced BFMs capable of interpreting and executing high-level goals, including language instructions and multi-task objectives. A notable example is MoConVQ [71], which introduces a unified motion control framework based on discrete latent codes learned via vector quantized variational autoencoder (VQ-VAE). The model supports a wide range of downstream tasks—including motion tracking, interactive control, and text-to-motion generation—by offering a compact and modular representation. MoConVQ also integrates with large language models (LLMs) and enables the simulated agents to be directed via in-context language prompts, thereby bridging symbolic reasoning and physical control.
Meanwhile, MaskedMimic [73] addresses physics-based character control as a general motion inpainting problem, producing full-body motions from partial descriptions like masked keyframes, objects, or text instructions. MaskedMimic involves a two-phase training process: first, a fully-constrained motion tracking controller learns to imitate diverse reference motions, then a partially-constrained VAE-based policy distills this knowledge through masked goal conditioning. As a result, MaskedMimic can dynamically adapt to complex scenes and support applications ranging from VR control to complex human-object interaction (HOI), as illustrated in Fig. 4. In addition, InterMimic [70] focuses on the HOI scenario and designs a two-stage teacher-student framework that distills imperfect motion capture interaction data into robust and physics-based controllers. Teacher policies are trained on subsets of noisy data and refined through simulation, then distilled into a student policy with RL-based fine-tuning. This curriculum strategy enables generalization across diverse interactions with high physical fidelity.
@@
Figure 4: MaskedMimic [73] is a representative BFM trained via goal-conditioned learning, which supports diverse control modalities and achieves seamless transitions between distinct tasks.
@@
For real-world robotic applications, HOVER [61] introduces a multi-mode policy distillation framework that allows humanoid robots to switch seamlessly between tasks like locomotion, manipulation, and navigation using a single unified policy distilled from an oracle. This eliminates the need for task-specific controllers, demonstrating general-purpose control in real-world environments, akin to the versatility seen in MaskedMimic for virtual characters.
@@
All the above methods follow the idea of BFM, which is revealed from two aspects: (i) their ability to learn a broad behavior coverage from diverse data sources, and (ii) their fast adaptation ability to downstream tasks. These models are trained on large-scale datasets and can be generalized across a wide range of motor skills, such as locomotion, HOI, and task-specific behaviors, without being limited to a single task. For instance, InterMimic is capable of handling various HOT tasks, while MoConVQ adapts to different tasks, such as goal-reaching and text-conditioned motion generation. Additionally, these models exhibit fast adaptation to new tasks, with minimal retraining, demonstrating their capacity to apply learned behaviors to new and unseen scenarios. Thus, the broad behavior coverage combined with the adaptation ability collectively characterizes these methods as BFMs.
3.1.2 Intrinsic Reward-driven Learning
@@
In tracking-based learning, the agent is consistently provided with an explicit objective (e.g., joint angles or velocities) and trained via a well-specified reward function to achieve targeted skill acquisition. In contrast, intrinsic reward-driven learning presents a distinct approach, where the agent is motivated to explore the environment without relying on explicit task-specific rewards. Instead, the agent is guided by intrinsic rewards, which are self-generated signals that encourage exploration, skill acquisition, or novelty detection. Extensive strategies for intrinsic reward-driven learning have been developed, including curiosity-driven exploration [75, 77, 90, 91], skill discovery [92, 76, 93, 78], and maximizing data coverage [94, 95, 80, 79], each encouraging the agent to explore different aspects of the environment. Denote by $I:\mathcal{S}\rightarrow\mathbb{R}$ the intrinsic reward function, the optimization objective in Eq. (1) is reformulated as
$J_{\pi}=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\big{(}r(\bm{s}_{t})+\beta_{t}\cdot I(\bm{s}_{t})\big{)}\right],$ (2)
@@
For example, [75] introduces an intrinsic curiosity module (ICM) that encourages the agent to explore unfamiliar states by providing an intrinsic reward based on the prediction error between the current and predicted next state. The intrinsic reward is proportional to the discrepancy between the predicted state and the actual state, motivating the agent to interact with environments that it cannot predict, thereby driving exploration and learning. ICM has been shown to significantly improve the agent’s ability to explore complex environments with sparse or no external rewards. In contrast, DIAYN [76] uses latent variable discovery to guide the agent’s exploration by encouraging the agent to maximize the diversity of the behaviors it exhibits. DIAYN introduces an intrinsic reward based on the mutual information between the agent’s latent skill variable and its environment state, encouraging the agent to discover and explore a wide range of distinct behaviors. By learning a set of diverse, reusable skills, DIAYN enables agents to tackle complex tasks without requiring domain-specific rewards, making it a valuable approach for unsupervised skill discovery.
@@
In addition, RE3 [80] focuses on state coverage maximization, where the intrinsic reward is driven by the state visitation frequency. The goal is to encourage the agent to explore states that are infrequent or underrepresented in its past experience. By learning from these underexplored regions, RE3 enhances the agent’s understanding of the environment and helps to avoid getting stuck in local optima. RE3’s approach allows the agent to explore a broader range of states, improving the diversity of the learned representations and making it suitable for environments with sparse external rewards. Following DIAYN and RE3, ODPP [81] is proposed as a unified framework to discover skills that are both diverse and have superior state coverage, based on a novel use of determinant point process. Specifically, the unsupervised objective is to maximize (i) the number of modes covered within each individual trajectory to enhance state coverage, and (ii) the number of modes across the trajectory space to promote skill diversity.
These unsupervised RL agents are considered BFMs based on two key observations: (i) they can effectively explore the environment and action space and discover generalizable behaviors under the motivation of intrinsic rewards, and (ii) they exhibit the capability to effectively learn and adapt to various tasks, as demonstrated in [79] and [96]. However, BFMs trained solely with intrinsic rewards face significant limitations. The agent often requires a huge amount of training to achieve broad behavior coverage under the guidance of intrinsic rewards, while consistently producing unreliable motion prior (e.g., unsafe or impractical motions), especially for humanoid robots with extremely complex dynamics [44]. Despite the convenience of this paradigm, future work may address these fundamental challenges by developing hybrid approaches that combine the exploratory benefits of intrinsic rewards with the task-relevant utility guarantees of goal-conditioned learning, achieving reliable BFMs for humanoid robots.
3.1.3 Forward-backward Representation Learning
@@
Recent advances in BFM are propelled by a novel framework entitled forward-backward (FB) representation learning [97] that disentangles the policy learning from task-specific objectives, demonstrating a fundamentally different approach with goal-conditioned learning and intrinsic reward-driven learning. By learning a universal policy representation, it can be rapidly adapted to new tasks through reward inference or demonstration alignment, without additional environment interaction or policy optimization.
@@
At its core, the FB representation learning seeks to learn a finite-rank approximation of the successor measure, which is an extension of the successor representation [98, 99]. It depicts the discounted future state visitation distribution as a measure over states. For each policy $\pi$, its successor measure is defined as
@@
The successor measure satisfies a measure-valued Bellman equation [97]:
@@
$=$ $\int_{\bm{s}^{\prime}\in\mathcal{S}}M^{\pi}\left(\mathrm{d}\bm{s}^{\prime}|\bm{s},\bm{a}\right)r\left(\bm{s}^{\prime}\right)$
@@
Notably, [82] proposes an estimation of the success measure as
$M^{\pi}(\mathcal{X}|\bm{s},\bm{a})\approx\int_{\bm{s}^{\prime}\in\mathcal{X}}F^{\pi}(\bm{s},\bm{a})^{\top}B(\bm{s}^{\prime})\rho(\mathrm{d}\bm{s}^{\prime}),$ (6)
where $\rho$ is an arbitrary distribution over states, $F^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ is the forward embedding and $B:\mathcal{S}\rightarrow\mathbb{R}^{d}$ is the backward embedding, respectively. Denote by $\bm{z}=\mathbb{E}_{\bm{s}\sim\rho}\left[B(\bm{s})r(\bm{s})\right]$, the action-value function is rewritten as
$Q^{\pi}_{r}=F^{\pi}(\bm{s},\bm{a})^{\top}\bm{z}.$ (7)
To learn a family of polices, [82] suggests that both the forward embedding $F^{\pi}$ and policy $\pi$ can be parameterized by the same task encoding vector $\bm{z}$, such that
$M^{\pi_{\bm{z}}}(\mathcal{X}|\bm{s},\bm{a})\approx\int_{\bm{s}^{\prime}\in\mathcal{X}}F^{\pi}(\bm{s},\bm{a},\bm{z})^{\top}B(\bm{s}^{\prime})\rho(\mathrm{d}\bm{s}^{\prime}),$ (8)
@@
As introduced in Section 1.3, [44] first introduced the term ”BFM” literally and proposed FB-IL for fast IL based on BFMs, which supports multiple IL principles, such as behavioral cloning, feature matching, and goal-based reduction, without needing separate RL routines for each new task. [45] further enhances the FB framework by incorporating auto-regressive features for more precise task encoding and better task representations in BFMs. While the standard FB method uses a linear task projection that can blur rewards and reduce spatial precision, auto-regressive features improve expressivity and performance, especially for tasks requiring spatial accuracy or generalization. Additionally, [45] introduces advantage-weighted regression (AWR) to address challenges with offline learning from complex datasets. The modified FB approach, FB-AWARE, combines auto-regressive features with advantage weighting, which performs well across new environments and even matches the performance of standard offline RL agents in benchmarks like D4RL.
Figure 7: Motivo [46] learns broad behavior priors and demonstrates outstanding zero-shot adaptation capability to diverse downstream tasks, including complex motion imitation, pose reaching, and composite reward optimization. Moreover, Motivo achieves real-time motor control while ensuring motion naturalness.
The FB framework provides a general and flexible approach for training BFMs by learning successor measure representations and applying pre-trained policies to new tasks. However, it suffers from several limitations: (i) when the latent dimension $d$ is finite, it relies on the assumption of a low-rank dynamic, leading to a limited inductive bias for policy selection; (ii) poor coverage in the training dataset causes offline learning to fail in reliably optimizing policies, often collapsing to a few suboptimal behaviors with weak performance on downstream tasks. These limitations greatly hinder the application of the FB framework for humanoid robots. To address these limitations, [46] proposes FB with conditional policy regularization (FB-CPR) and introduces Motivo, as illustrated in Fig. 7, which solves diverse humanoid WBC tasks in a zero-shot manner, including motion tracking, goal reaching, and reward optimization.
@@
$L_{\rm FB-CPR}=$ $-\mathbb{E}_{\bm{z}\sim\nu,\bm{s}\sim\mathcal{D}_{\rm online},\bm{a}\sim\pi_{\bm{z}}(\cdot|\bm{s})}\left[F(\bm{s},\bm{a},\bm{z})^{\top}\bm{z}\right]$ (13)
@@
where $\mathcal{D}_{\rm online}$ is the associated replay buffer of unsupervised transitions, $p_{\pi}(\bm{s},\bm{z})$ is the joint distribution of $(\bm{s},\bm{z})$ induced by FB, and $p_{\mathcal{E}}$ is the joint distribution of the dataset. However, it is intractable to optimize the divergence term directly via a RL procedure. To tackle the problem, FB-CPR interprets the divergence as an expected return under the policies and defines a divergence-based reward $r_{\rm div}$ as
@@
Then, a discriminator network $D:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is trained to estimate the $r_{\rm div}$:
$r_{\rm div}=\log\frac{p_{\pi}(\bm{s},\bm{z})}{p_{\mathcal{E}}(\bm{s},\bm{z})}\approx\log\frac{D}{1-D}.$ (15)

### v4

@@ III-A1 Goal-conditioned Learning
TeamPlay [liu2022motor], ASE [peng2022ase], PHC [luo2023perpetual], CALM [tessler2023calm], MTM [wu2023masked],
CASE [dou2023case], InterMimic [xu2025intermimic], MoConVQ [yao2024moconvq], H-GAP [jiang2024hgap],
MaskedMimic [tessler2024maskedmimic], HugWBC [xue2025unified], HOVER [he2025hover], ModSkill [huang2025modskill],
CLONE [li2025clone], TWIST [ze2025twist], TWIST2 [ze2025twist2], AMS [pan2025agility],
Any2Track [zhang2025track], BFM4Humanoid [zeng2025behavior], SONIC [luo2025sonic]
@@
ICM [pathak2017curiosity], DIAYN [eysenbach2018diversity], RND [burda2018exploration], APS [liu2021aps],
ProtoRL [yarats2021reinforcement], RE3 [seo2021state], ODPP [chen2023unified]
@@
FB [touati2021learning], FB-IL [pirotta2024fast], FB-AW [cetin2025finer], FB-ARE [cetin2025finer],
FB-AWARE [cetin2025finer], Motivo [tirinzoni2025zeroshot], BFM-Zero [li2025bfm]
Adaptation
Adaptation
@@
TokenHSI [pan2025tokenhsi], Belief-FB [bobrin2025zero], Rotation-FB [bobrin2025zero],
TaskTokens [vainshtein2025task], ReLA [sikchi2025fast], LoLA [sikchi2025fast]
@@
UniHSI [xiao2024unified], TokenHSI [pan2025tokenhsi], UniPhys [wu2025uniphys], CLoSD [tevet2025closing],
LangWBC [shao2025langwbc], LeVERB [xue2025leverb], BeyondMimic [liao2025beyondmimic], SENTINEL [wang2025sentinel]
Skill learning from motion tracking. Among the diverse approaches to goal-conditioned learning, tracking-based learning represents a specialized form where the target behavior is explicitly defined by dense reference supervision or guidance, typically derived from motion capture data or expert demonstrations. At each time step, the agent is often trained to track the given reference motion’s joint angle or kinematic pose of the next time step [peng2018deepmimic]. The primary motivation behind tracking-based learning is that learning to track a single pose is more achievable and general than directly imitating a whole motion, especially a complex motion.
For example, the TeamPlay [liu2022motor] framework trains an agent to imitate a large amount of football motion capture data via a DeepMimic-like [peng2018deepmimic] approach, aiming to learn a complete behavior set required for the football game. Then the agent is leveraged to sample substantial state-action pairs to train a neural probabilistic motor primitive (NPMP) model [merel2018neural] and derive a low-level latent-conditioned controller. Finally, the controller is applied for further drill learning by conducting RL with a drill-specific (e.g., follow, dribble, shoot, and kick-to-target) reward function. Here, the learned low-level controller can be viewed as a BFM as it learns realistic human-like movements based on motion capture data and can be rapidly adapted to diverse higher-level drill learning.
Similarly, the adversarial skill embeddings (ASE) framework [peng2022ase] learns a reusable latent space of motor skills by combining adversarial IL with unsupervised RL. Trained on unstructured motion data, ASE produces a latent-conditioned low-level controller capable of generating diverse and physically plausible behaviors, serving as a general-purpose motor prior to downstream tasks. Building on ASE, the conditional adversarial latent models (CALM) framework [tessler2023calm] incorporates a conditional discriminator to enable fine-grained control over generated motions through latent manipulation. This line of work is further extended by CASE [dou2023case], which introduces skill-conditioned IL with training techniques such as focal skill sampling and skeletal residual forces to enhance agility and motion diversity.
@@
While the methods above achieve efficient skill acquisition from large behavior datasets, the HugWBC [xue2025unified] method explores learning versatile locomotion skills without relying on pre-collected motion data. The framework automatically generates adaptive behaviors through a structured RL process, where a general command space dynamically produces feasible velocity, gait, and posture targets during training. By reformulating WBC as a self-supervised command-tracking problem, the work establishes a new direction for developing general-purpose humanoid controllers that learn robust skills through environmental interaction rather than data imitation.
@@
Moving beyond body-level skill learning, ModSkill [huang2025modskill] introduces a modular framework that decouples full-body motion into part-specific skills for individual body parts. This modularization allows for efficient and scalable learning, as each body part is controlled independently by a low-level controller driven by part-specific skill embeddings. ModSkill’s ability to focus on body-part-level skills makes it a powerful system for controlling complex motions and adapting learned behaviors across different tasks. By utilizing a skill modularization attention layer, ModSkill enhances the generalization of motor skills across various tasks, such as reaching or striking, thereby further improving task-specific adaptation.
@@
From primitive skills to high-level goal execution. The success of BFMs in learning diverse primitive skills has propelled the development of more advanced BFMs capable of interpreting and executing high-level goals, including language instructions and multi-task objectives. A notable example is MoConVQ [yao2024moconvq], which introduces a unified motion control framework based on discrete latent codes learned via vector quantized variational autoencoder (VQ-VAE). The model supports a wide range of downstream tasks, including motion tracking, interactive control, and text-to-motion generation, by offering a compact and modular representation. MoConVQ also integrates with large language models (LLMs) and enables the simulated agents to be directed via in-context language prompts, thereby bridging symbolic reasoning and physical control.
@@
Figure 5: MaskedMimic [tessler2024maskedmimic] is a representative BFM trained via goal-conditioned learning, which supports diverse control modalities and achieves seamless transitions between distinct tasks.
Meanwhile, MaskedMimic [tessler2024maskedmimic] addresses physics-based character control as a general motion inpainting problem, producing full-body motions from partial descriptio, such as masked keyframes, objects, or text instructions. MaskedMimic involves a two-phase training process. First, a fully-constrained motion tracking controller is trained to imitate diverse reference motions. Then, a partially-constrained VAE-based policy distills this knowledge through masked goal conditioning. As a result, MaskedMimic can dynamically adapt to complex scenes and support applications ranging from VR control to complex human-object interaction (HOI), as illustrated in Fig. 5. In addition, InterMimic [xu2025intermimic] focuses on the HOI scenario and designs a two-stage teacher-student framework that distills imperfect motion capture interaction data into robust and physics-based controllers. Teacher policies are trained on subsets of noisy data and refined through simulation, then distilled into a student policy with RL-based fine-tuning. This curriculum strategy enables generalization across diverse interactions with high physical fidelity.
@@
For real-world robotic applications, HOVER [he2025hover] introduces a multi-mode policy distillation framework that enables humanoid robots to switch seamlessly between tasks (e.g., locomotion, manipulation, and navigation) using a single unified policy distilled from an oracle. This eliminates the need for task-specific controllers, demonstrating general-purpose control in real-world environments, akin to the versatility seen in MaskedMimic for virtual characters.
@@
All the above methods follow the idea of BFM, which is revealed from two aspects: (i) their ability to learn broad behavioral priors from diverse data sources, and (ii) their fast adaptation ability to downstream tasks. Trained on large-scale data, these models generalize across a wide range of behaviors without being limited to a single domain. For example, InterMimic handles various HOI tasks, while MoConVQ adapts to tasks like goal-reaching and text-conditioned motion generation. Critically, they exhibit rapid adaptation to new scenarios with minimal retraining. Thus, the learned broad behavioral priors, along with the swift adaptation capability, characterize these methods as BFMs.
III-A2 Intrinsic Reward-driven Learning
@@
In tracking-based learning, the agent is consistently provided with an explicit objective (e.g., joint angles or velocities) and trained via a well-specified reward function to achieve targeted skill acquisition. In contrast, intrinsic reward-driven learning presents a distinct approach, where the agent is motivated to explore the environment without relying on explicit task-specific rewards. Instead, as illustrated in Fig. 6, the agent is guided by intrinsic rewards, which are self-generated signals that encourage exploration, skill acquisition, or novelty detection. Extensive strategies for intrinsic reward-driven learning have been developed, including curiosity-driven exploration [pathak2017curiosity, burda2018exploration, pathak2019self, sekar2020planning], skill discovery [gregor2017variational, eysenbach2018diversity, hansen2020fast, liu2021aps], and maximizing data coverage [bellemare2016unifying, ostrovski2017count, seo2021state, yarats2021reinforcement], each encouraging the agent to explore different aspects of the environment. Denote by $r^{\rm int}:\mathcal{S}\rightarrow\mathbb{R}$ the intrinsic reward function, the optimization objective in Eq. (1) is reformulated as
$J_{\pi}=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\big(r^{\rm ext}(\bm{s}_{t})+\beta_{t}\cdot r^{\rm int}(\bm{s}_{t})\big)\right],$ (2)
@@
For example, the intrinsic curiosity module (ICM) [pathak2017curiosity] encourages the agent to explore unfamiliar states by providing an intrinsic reward based on the prediction error between the current and predicted next state. The intrinsic reward is proportional to the discrepancy between the predicted state and the actual state, motivating the agent to interact with environments that it cannot predict, thereby driving exploration and learning. ICM has been shown to significantly improve the agent’s ability to explore complex environments with sparse or no external rewards. In contrast, DIAYN [eysenbach2018diversity] utilizes latent variable discovery to guide the agent’s exploration, encouraging it to maximize the diversity of behaviors exhibited. DIAYN introduces an intrinsic reward based on the mutual information between the agent’s latent skill variable and its environment state, encouraging the agent to discover and explore a wide range of distinct behaviors. By learning a set of diverse, reusable skills, DIAYN enables agents to tackle complex tasks without requiring domain-specific rewards, making it a valuable approach for unsupervised skill discovery.
@@
In addition, RE3 [seo2021state] focuses on state coverage maximization, where the intrinsic reward is driven by the state visitation frequency. The goal is to encourage the agent to explore states that are infrequent or underrepresented in its past experience. By learning from these underexplored regions, RE3 enhances the agent’s understanding of the environment and helps prevent the agent from getting stuck in local optima. RE3 allows the agent to explore a broader range of states, improving the diversity of the learned representations and making it suitable for environments with sparse external rewards. Following DIAYN and RE3, ODPP [chen2023unified] is proposed as a unified framework to discover skills that are both diverse and have superior state coverage, based on a novel use of the determinant point process. Specifically, the unsupervised objective is to maximize (i) the number of modes covered within each trajectory to enhance state coverage, and (ii) the number of modes across the trajectory space to promote skill diversity.
Compared to BFMs trained on large-scale and diverse behavior datasets, intrinsic reward-driven models are typically trained on a single environment/task, which limits their ability to provide strong and intuitive behavioral priors. However, these models exhibit several key properties desirable for BFM pre-training. Specifically, they can (i) effectively explore the environment and action space and discover generalizable and latent behavioral priors under the motivation of intrinsic rewards, and (ii) demonstrate the capability to learn and adapt to various downstream tasks, as shown in [yarats2021reinforcement, laskin2urlb]. Therefore, we consider these methods as influential and conceptually related prototypes, whose unsupervised behavioral prior acquisition mechanisms can inspire the development of more effective BFMs. Future work may combine the exploratory advantages of intrinsic rewards with the task-specific utility of goal-conditioned learning, leading to a more comprehensive approach to behavioral prior learning.
III-A3 Forward-backward Representation Learning
@@
Recent advances in BFM are propelled by a novel framework entitled forward-backward (FB) representation learning [blier2021learning]. As illustrated in Fig. 7, it disentangles policy learning from task-specific objectives, which is fundamentally different from goal-conditioned learning and intrinsic reward-driven learning. By learning a universal policy representation, it can be rapidly adapted to new tasks through reward inference or demonstration alignment, without additional environment interaction or policy optimization.
@@
At its core, the FB representation learning seeks to learn a finite-rank approximation of the successor measure, which is an extension of the successor representation [dayan1993improving, kulkarni2016deep]. It depicts the discounted future state visitation distribution as a measure over states. For each policy $\pi$, its successor measure is defined as
@@
The successor measure satisfies a measure-valued Bellman equation [blier2021learning]:
@@
$=$ $\int_{\bm{s}^{\prime}\in\mathcal{S}}M^{\pi}\left(\mathrm{d}\bm{s}^{\prime}|\bm{s},\bm{a}\right)r\left(\bm{s}^{\prime}\right).$
@@
Notably, the success measure can be estimated as [touati2021learning]
$M^{\pi}(\mathcal{X}|\bm{s},\bm{a})\approx\int_{\bm{s}^{\prime}\in\mathcal{X}}F(\bm{s},\bm{a})^{\top}B(\bm{s}^{\prime})\rho(\mathrm{d}\bm{s}^{\prime}),$ (6)
where $\rho$ is an arbitrary distribution over states, $F:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$ is the forward embedding and $B:\mathcal{S}\rightarrow\mathbb{R}^{d}$ is the backward embedding, respectively. Denote by $\bm{z}=\mathbb{E}_{\bm{s}\sim\rho}\left[B(\bm{s})r(\bm{s})\right]$, the action-value function is rewritten as
$Q^{\pi}_{r}=F(\bm{s},\bm{a})^{\top}\bm{z}.$ (7)
To learn a family of polices, both the forward embedding $F$ and policy $\pi$ can be parameterized by the same task encoding vector $\bm{z}$ [touati2021learning], such that
$M^{\pi_{\bm{z}}}(\mathcal{X}|\bm{s},\bm{a})\approx\int_{\bm{s}^{\prime}\in\mathcal{X}}F(\bm{s},\bm{a},\bm{z})^{\top}B(\bm{s}^{\prime})\rho(\mathrm{d}\bm{s}^{\prime}),$ (8)
@@
As introduced in Section I-C, the term ”BFM” is first literally introduced in the FB-IL framework [pirotta2024fast], which supports multiple IL principles without needing separate RL routines for each new task. FB-AW [cetin2025finer] enhances the FB framework by incorporating the advantage-weighted (AW) regression loss in offline RL algorithms, achieving stable performance on complex datasets (e.g., DMC Humanoid [tassa2018deepmind]) composed of near-optimal trajectories. The subsequent FB-AWARE [cetin2025finer] method further employs an auto-regressive encoding (ARE) mechanism to let fine-grained task features depend on coarser-grained task information. While the standard FB method employs a linear task projection that can blur rewards and reduce spatial precision, this mechanism can significantly enhance the expressivity and performance of the FB framework, particularly for tasks that require spatial accuracy or generalization. Notably, the AW and ARE mechanisms can be flexibly combined based on the task requirements. For example, the ARE mechanism can be used independently in the FB-ARE variant for environments where the AW component is not required for satisfactory performance.
Figure 8: Motivo [tirinzoni2025zeroshot] learns broad behavior priors and demonstrates outstanding zero-shot adaptation capability to diverse downstream tasks, including complex motion imitation, pose reaching, and composite reward optimization. Moreover, Motivo achieves real-time motor control while ensuring motion naturalness.
The FB framework provides a general and flexible approach for training BFMs by learning successor measure representations and applying pre-trained policies to new tasks. However, it suffers from several limitations: (i) when the latent dimension $d$ is finite, it relies on the assumption of a low-rank dynamic, leading to a limited inductive bias for policy selection; (ii) poor coverage in the training dataset causes offline learning to fail in reliably optimizing policies, often collapsing to a few suboptimal behaviors with weak performance on downstream tasks. These limitations greatly hinder the application of the FB framework for humanoid robots. To address these limitations, the Motivo framework [tirinzoni2025zeroshot] combines the FB framework with a conditional policy regularization (FB-CPR) term. As illustrated in Fig. 8, Motivo solves diverse humanoid WBC tasks in a zero-shot manner, including motion tracking, goal reaching, and reward optimization.
@@
$L_{\rm FB-CPR}=$ $-\mathbb{E}_{\begin{subarray}{c}\bm{z}\sim\nu,\\ \bm{s}\sim\mathcal{D}_{\rm online},\bm{a}\sim\pi_{\bm{z}}(\cdot|\bm{s})\end{subarray}}\left[F(\bm{s},\bm{a},\bm{z})^{\top}\bm{z}\right]$ (13)
@@
where $\mathcal{D}_{\rm online}$ is the associated replay buffer of unsupervised transitions, $p_{\pi}(\bm{s},\bm{z})$ is the joint distribution of $(\bm{s},\bm{z})$ induced by FB, and $p_{\mathcal{E}}$ is the joint distribution of the dataset. However, it is intractable to optimize the divergence term directly via a RL procedure. To tackle the problem, FB-CPR interprets the divergence as an expected return under the policies and defines a divergence-based reward $r^{\rm div}$ as
@@
Then, a discriminator network $D:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is trained to estimate the $r^{\rm div}$:
$r^{\rm div}=\log\frac{p_{\pi}(\bm{s},\bm{z})}{p_{\mathcal{E}}(\bm{s},\bm{z})}\approx\log\frac{D}{1-D}.$ (15)